OpenStudy (aravindg):

Note :This is not a question ! Vedic Mathematics tutorials (contd...) ...Hope you find it useful ! Pleaase provide your feedback !happy studying !

OpenStudy (aravindg):

\[\huge VEDIC \;\;\;\; MATHS \]

OpenStudy (aravindg):

.Adding Time _______________ Here is a nice simple way to add hours and minutes together: Let's add 1 hr and 35 minutes and 3 hr 55 minutes together. What you do is this: make the 1 hr 35 minutes into one number, which will give us 135 and do the same for the other number, 3 hours 55 minutes, giving us 355 Now you want to add these two numbers together: 135 355 ___ 490 So we now have a sub total of 490. What you need to do to this and all sub totals is add the time constant of 40. No matter what the hours and minutes are, just add the 40 time constant to the sub total. 490 + 40 = 530 So we can now see our answer is 5 hrs and 30 minutes!

OpenStudy (aravindg):

Multiplication of 99 with any number (You can't use this rule for other numbers) Let me explain this rule by taking examples ____________________________________________ 1. 46*99 = 46*(100-1) = 46*100-46 = 4600-46 = 4554. 2. 362*99 = 362*(100-1) = 36200-362 = 35838. Understand the rule by 1 more example 2841*99 = 2841*(100-1) = 284100-2841 = 281259

OpenStudy (aravindg):

Square of a 2 digit number ending with 5 (You can't use this rule for other numbers) _______________________________________________________________________________ Let me explain this rule by taking examples 35^2 = 3*(3+1) ---25 = 1225 Example. Other example 95^2 = 9*(9+1) ---25 = 9025

OpenStudy (aravindg):

Multiplication of any two numbers, both ranging from 11 to 19. (You can’t use this rule for other numbers) ____________________________________________________________________________________ Let me explain this rule by taking examples 13*19 = (13+9)*10 + (3*9) = 220 + 27 = 247 Means add first number and last digit of the second number take zero in the third place of this number then add product of last digit of the two numbers in it. Example. 18*14 = (18+4)*10 + (8*4) = 220 + 32 = 252

OpenStudy (aravindg):

wel thats end of today's tutorial :) hope you all found it useful ! P.S. wish i was better at latexing could have made it more attractive :)

OpenStudy (anonymous):

NICE.............!!!!!!!!

OpenStudy (aravindg):

thanks @nitz

OpenStudy (hba):

Have you given multiplication of 11 with any number ?

OpenStudy (mathmate):

@AravindG Nice! Post others soon.

OpenStudy (aravindg):

@hba yep it was there in my previous tutorial

OpenStudy (aravindg):

@mathmate thanks :)

OpenStudy (mayankdevnani):

good job...... @AravindG

OpenStudy (aravindg):

thanks @mayankdevnani

OpenStudy (mayankdevnani):

can you solve it and then clear to me. 29*15=? in 10 sec.

OpenStudy (mayankdevnani):

@AravindG can you???

OpenStudy (aravindg):

(30-1)(30)/2

OpenStudy (mayankdevnani):

plz reply by drawing!!!!

OpenStudy (aravindg):

900-30/2

OpenStudy (aravindg):